The expression (log0.52 8)1/2 has the value equal to ______.

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
given (log0.52 8)1/2=1/4 log0.5 8 from log property (loganx)m=m/n(loga x)
also 1/4(log0.58)= 1/4((log10 8)/log10 0.5)=1/4((log 23)/log 2-1)=-3/4